# DexGarmentLab: Dexterous Garment Manipulation Environment with Generalizable Policy

### v2

@@ Appendix G DexGarmentLab Assets
Appendix H Detailed Task Description
@@
Fold Tops (Sec. H.1), Fold Dress (Sec. H.2), Fold Trousers (Sec. H.3), Fling Tops (Sec. H.4), Fling Dress (Sec. H.5), Fling Trousers (Sec. H.6).
@@
Hang Coat (Sec. H.7), Hang Tops (Sec. H.8), Hang Dress (Sec. H.9), Hang Trousers (Sec. H.10), Store Tops (Sec. H.11), Wear Baseball Cap (Sec. H.12), Wear Bowl Hat (Sec. H.13), Wear Scarf (Sec. H.14), Wear Glove (Sec. H.15).
@@ H.1 Fold Tops
H.1.1 Task Initialization and Randomization
@@ H.1.2 Task Sequence
H.1.3 Task Success Metrics
@@
H.1.4 Garment Assets for Task
@@ H.2 Fold Dress
H.2.1 Task Initialization and Randomization
@@ H.2.2 Task Sequence
H.2.3 Task Success Metrics
@@
H.2.4 Garment Assets for Task
@@ H.3 Fold Trousers
H.3.1 Task Initialization and Randomization
@@ H.3.2 Task Sequence
H.3.3 Task Success Metrics
@@
H.3.4 Garment Assets for Task
@@ H.4 Fling Tops
H.4.1 Task Initialization and Randomization
@@ H.4.2 Task Sequence
H.4.3 Task Success Metrics
@@
H.4.4 Garment Assets for Task
@@ H.5 Fling Dress
H.5.1 Task Initialization and Randomization
@@ H.5.2 Task Sequence
H.5.3 Task Success Metrics
@@
H.5.4 Garment Assets for Task
@@ H.6 Fling Trousers
H.6.1 Task Initialization and Randomization
@@ H.6.2 Task Sequence
H.6.3 Task Success Metrics
@@
H.6.4 Garment Assets for Task
@@ H.7 Hang Coat
H.7.1 Task Initialization and Randomization
@@ H.7.2 Task Sequence
H.7.3 Task Success Metrics
@@
H.7.4 Garment Assets for Task
@@ H.8 Hang Tops
H.8.1 Task Initialization and Randomization
@@ H.8.2 Task Sequence
H.8.3 Task Success Metrics
@@
H.8.4 Garment Assets for Task
@@ H.9 Hang Dress
H.9.1 Task Initialization and Randomization
@@ H.9.2 Task Sequence
H.9.3 Task Success Metrics
@@
H.9.4 Garment Assets for Task
@@ H.10 Hang Trousers
H.10.1 Task Initialization and Randomization
@@ H.10.2 Task Sequence
H.10.3 Task Success Metrics
@@
H.10.4 Garment Assets for Task
@@ H.11 Store Tops
H.11.1 Task Initialization and Randomization
@@ H.11.2 Task Sequence
H.11.3 Task Success Metrics
@@
H.11.4 Garment Assets for Task
@@ H.12 Wear Baseball Cap
H.12.1 Task Initialization and Randomization
@@ H.12.2 Task Sequence
H.12.3 Task Success Metrics
@@
H.12.4 Garment Assets for Task
@@ H.13 Wear Bowl Hat
H.13.1 Task Initialization and Randomization
@@ H.13.2 Task Sequence
H.13.3 Task Success Metrics
@@
H.13.4 Garment Assets for Task
@@ H.14 Wear Scarf
H.14.1 Task Initialization and Randomization
@@ H.14.2 Task Sequence
H.14.3 Task Success Metrics
@@
H.14.4 Garment Assets for Task
@@ H.15 Wear Glove
H.16 Additional Explanation for Success Metric
@@
Appendix I Broader Impact

### v3

@@ I.3 Generalizable Policy Limitation
Appendix K Detailed Task Description
@@
Fold Tops (Sec. K.1), Fold Dress (Sec. K.2), Fold Trousers (Sec. K.3), Fling Tops (Sec. K.4), Fling Dress (Sec. K.5), Fling Trousers (Sec. K.6).
@@
Hang Coat (Sec. K.7), Hang Tops (Sec. K.8), Hang Dress (Sec. K.9), Hang Trousers (Sec. K.10), Store Tops (Sec. K.11), Wear Baseball Cap (Sec. K.12), Wear Bowl Hat (Sec. K.13), Wear Scarf (Sec. K.14), Wear Glove (Sec. K.15).
@@
K.1.1 Task Initialization and Randomization
@@
K.1.3 Task Success Metrics
@@
K.1.4 Garment Assets for Task
@@
K.2.1 Task Initialization and Randomization
@@
K.2.3 Task Success Metrics
@@
K.2.4 Garment Assets for Task
@@
K.3.1 Task Initialization and Randomization
@@
K.3.3 Task Success Metrics
@@
K.3.4 Garment Assets for Task
@@
K.4.1 Task Initialization and Randomization
@@
K.4.3 Task Success Metrics
@@
K.4.4 Garment Assets for Task
@@
K.5.1 Task Initialization and Randomization
@@
K.5.3 Task Success Metrics
@@
K.5.4 Garment Assets for Task
@@
K.6.1 Task Initialization and Randomization
@@
K.6.3 Task Success Metrics
@@
K.6.4 Garment Assets for Task
@@
K.7.1 Task Initialization and Randomization
@@
K.7.3 Task Success Metrics
@@
K.7.4 Garment Assets for Task
@@
K.8.1 Task Initialization and Randomization
@@
K.8.3 Task Success Metrics
@@
K.8.4 Garment Assets for Task
@@
K.9.1 Task Initialization and Randomization
@@
K.9.3 Task Success Metrics
@@
K.9.4 Garment Assets for Task
@@
K.10.1 Task Initialization and Randomization
@@
K.10.3 Task Success Metrics
@@
K.10.4 Garment Assets for Task
@@
K.11.1 Task Initialization and Randomization
@@
K.11.3 Task Success Metrics
@@
K.11.4 Garment Assets for Task
@@
K.12.1 Task Initialization and Randomization
@@
K.12.3 Task Success Metrics
@@
K.12.4 Garment Assets for Task
@@
K.13.1 Task Initialization and Randomization
@@
K.13.3 Task Success Metrics
@@
K.13.4 Garment Assets for Task
@@
K.14.1 Task Initialization and Randomization
@@
K.14.3 Task Success Metrics
@@
K.14.4 Garment Assets for Task
@@
K.16 Additional Explanation for Success Metric
@@
Appendix L Broader Impact